the surface of earth is 71% covered by water according to the USGS, there is 24,064,000 km^3 of water in ice caps, glaciers and snow. If all the ice and snow were to melt and flow into the oceans, calculate how much the average sea level rise.
Answer #1

Calculation of Sea Level Rise from Melting Ice

Given:

  • Total ice volume = 24,064,000km3=2.4064×107km3

  • Earth’s surface area = 510.1×106km2

  • Ocean coverage = 71% → Ocean area = 0.71×510.1×106=3.6217×108km2

Formula:

Sea level rise=Volume of melted iceOcean surface area

Calculation:

Sea level rise=2.4064×107km33.6217×108km2=0.0664km=66.4m

Answer:
If all ice melted, the average sea level would rise by ~66.4 meters.


answered by: anonymous